OpenStudy (anonymous):

Marylou is blowing up a beach ball. The package says that the ball has a diameter of 18 inches. What is the volume of the beach ball when it is completely full? Use 3.14 for . A, 24,416.6400000000 cubic inches B, 3,052.08 cubic inches C. 1,017.36 cubic inches D. 339.12 cubic inches @IGREEN

OpenStudy (igreen):

\(\sf V = \dfrac{4}{3} \pi r^3\) Plug in what we know: \(\sf V = \dfrac{4}{3}(3.14)(9^3)\) Simplify

OpenStudy (anonymous):

so its C ? @IGREEN

OpenStudy (igreen):

No..check again. \(\sf \dfrac{4}{3}(3.14)(9^3) \rightarrow \dfrac{4}{3} \times 3.14 \times 9 \times 9 \times 9\)

OpenStudy (anonymous):

sorry about that its B @IGREEN

geerky42 (geerky42):

Yep.

OpenStudy (anonymous):

Thank you could you help me with a few more ? @geerky42